# Actminer: Applying Causality Tracking and Increment Aligning for Graph-based Threat Hunting

### v1

@@ IV System Design
IV-A System Overview
Data Preparation Module (§ IV-B). The attack query graphs are extracted from threat intelligence reports, and provenance graphs are constructed based on extensive underlying logs. Duplicate events and orphan nodes within the provenance graphs undergo filtering, which is a necessary and common practice in existing work [32].
Casual Relation and Semantic Processing Module (§ IV-C). When a new attack query graph or provenance graph generated, Actminer will first categorize the entities into four classes. Then the provenance graph will delivery to the next module while the attack query graph still need to be processed. Next, Actminer merges analogous actions in the attack query graph. Finally, Actminer employs Equivalent Semantic Transfer which traces potentially overlooked attack chains by tracking malicious semantics, to identify the suspicious actions in the next module.
Threat Hunting and Incremental Aligning Module (§ IV-D). Actminer will hunt attack-related scenario by chaining suspicious semantic nodes and generating suspicious semantic tree. As time progresses, batch data is continuously inputted into the Actminer, persistently updating our suspicious semantic trees and unveiling more latent malicious attacks. And to control the memory consumption, Actminer will store the unupdated tree branch to the database unless certain behavior related to this branch.
The basic architecture of Actminer is shown in Figure 2, which can be divided into three modules: (I) the Data Preparation Module, (II) Casual Relation and Semantic Processing Module, and (III) Threat Hunting and Incremental Aligning Module. It is important to note that Actminer continuously runs the above three modules as the time progresses. Details of system design for each module are given in Section IV-B, Section IV-C, Section IV-D, respectively.
@@ IV-A System Overview
IV-B Data Preparation Module
@@
IV-B1 Provenance Preparation
Provenance graphs are composed of log data collected from various platforms by data collectors. In this work, we employ open-source tools such as eAuditd [33], Kellect [34], and Event Tracing for Windows (ETW) [35] to gather relevant system logs from Linux and Windows environments. Actminer transforms each event into a directed, time-stamped edge, in which the source node represents the object being acted upon. For any event $et\in E$, Actminer represents it as a quintuple $\langle{UID_{s},UID_{o},OP,Ti_{i}}\rangle$. $UID_{s}$ and $UID_{o}$ are unique identifiers for the subject and object of $et$, respectively. $OP$ denotes the type of $et$, and $Ti_{i}$ denotes the time when $et$ occurred.
Directly processing such massive raw log data is extremely challenging. To address this, we perform pruning operations on the low-level log data. Specifically, redundant events without context are removed [36, 9]. This means if the subject UID ($UID_{s}$), object UID ($UID_{o}$), and operation ($OP$) are identical, and the timestamps ($T_{i}$) are consecutive, the most recent timestamp will be preserved. Furthermore, our methodology includes the removal of isolated nodes within the provenance graph. The isolated nodes refer to entities that lack any incoming or outgoing edges. For example, we find that the data contains many these nodes manifest without any parent or child nodes, none of the events contain a subject UID ($UID_{s}$) or object UID ($UID_{o}$) matching the UID of the node, lack contextual information and fail to provide meaningful insights. So remove them do not compromise the integrity of the graph representation.
Concurrently, during the process of constructing provenance graphs, the node types are categorized based on the entity type contained within the logs. For instance, $F_{a}$ represents files involving user and system-sensitive information, such as the boot.ini file on Windows and /etc/passwd file on Linux. The specific categorization is detailed in Table I.
As show in Table I, we are inspired by APTShield [9], Conan [8] and refined the classification heuristics optimized the classification method we obtained from POIROT. Furthermore, we extensively gather CTI reports from various online sources and network channels, identifying specific file paths that exhibit heightened susceptibility to attacks. Consequently, we adapt the importance degree of these paths based on their frequency of occurrence to finally obtain ten distinct labels.
TABLE I: A categorization of distinct entities and their corresponding label assignments.
@@ IV-B1 Provenance Preparation
By analyzing the CDM18 and CDM19 which refer to the data definitions for DARPA’s E3 and E4 programs, respectively, we adopt events based on a few general fields in the CDM (i.e., the events of read, write, fork, clone, create, execute, load, and inject), which were most commonly used in previous studies such as Holme [3], Sleuth [10], and Morse [4]. The details are show in Table II.
TABLE II: Classification of different event types.
@@
IV-B2 Query Graph Preparation.
@@
CTI reports describe attacks that have already occurred. We collect the latest threat intelligence from websites such as Microsoft and Symantec. Leveraging open-source tools like Extractor [28], we extract attack query graphs ($G_{q}$) from cyber threat intelligence (CTI) reports. Analogous to provenance graphs, upon extracting attack query graphs, the constituent entities undergo a corresponding mapping process.
IV-C Casual Relation and Semantic Processing Module
@@
In this section, we sequentially describe the process of Module II in Figure 2, i.e., merging analogous actions and equivalent semantic transfer. In a nutshell, analogous actions in attack query graphs are merged while employing an equivalent semantic transfer strategy. This process enhances attack hunting capabilities while simultaneously preparing the data for input to Module III.
IV-C1 Merging Analogous Actions
@@
When Actminer adopts the attack query graphs to perform the threat hunting tasks, it encounters the issue of rigid matching. This situation leads the alignment between attack query graphs and provenance graphs become hard to achieve. Finally, it would result in high false negative and positive rate in the Module III.
@@
Additionally, the directly extracted $G_{q}$ is fixed, which limits its ability to generalize and effectively defend against variants of known attacks, as it is specifically tailored to documented attack patterns. To address this challenge, there is a key insights of Actminer: merging analogous actions. This method integrates nodes with similar operations. For instance, instead of representing each individual file with a distinct node, the generalized $G_{q}$ may group files based on their types (e.g., $F_{a},F_{b},F_{c}$ in Table I) in the attack chain. Simultaneously, we consider the temporal relationships of attack events. Processes can be merged based on their functional attributes or the operations they perform, rather than being strictly tied to specific executable paths or process names. In essence, this method is to transform the information flow from the source entity to the same target entity while preserving the semantic meaning of the source entity, so that equivalent events can be removed as redundant information, thereby achieving efficient compression of the data.
IV-C2 Equivalent Semantic Transfer
@@
During the attack process, malicious processes controlled by the attacker interact with other entities [8], causing malicious behaviors and effects to proliferate across the system through the intricate web of entity interactions and information propagation pathways, thereby expanding the attacker’s control scope.
@@
Based on above finding, we construct the equivalent semantics transitivity strategy to address the false negative issue. The key insight of this strategy is that the semantics of malicious behavior propagate with taking actions. As shown in Figure 3.(a), it depicts a short attack path described in the query graph, where a process controlled by the attacker tampers with sensitive files. Figure 3.(b) illustrates one of the attacker’s specific implementation approaches: First, a process P1 with suspicious semantics injects malicious code to create process P2; then, process P2 writes a suspicious file Fe; next, another malicious process P3 accesses the tainted file; subsequently, the process P3 with suspicious semantics creates process P4; finally, process P4 with suspicious semantics performs a write operation on sensitive files. Through this series of operations, the ultimate behavior of a suspicious process tampering with sensitive files Fa is achieved. Although the types of controlled entities change during the attack process, the ultimate goal remains unchanged. The operations in (b) are essentially the same as (a) and align with the attacker’s intent, making the two paths semantically equivalent.
@@
To comprehensively capture attack intents and mitigate attack evasion while addressing semantic gaps between $G_{q}$ and $G_{p}$, we integrated log data from the DARPA project, leveraged attack stage theories proposed in previous works like Conan [8] and Aptshield [9], and incorporated relevant descriptions from the CDM document [37] to extract six equivalent semantic transitivity policies, as illustrated in Table III. Therefore, Actminer can track and locate suspicious behavior on the host in real-time by analyzing the data flow transmission, while systems like HOLMES [3] cannot do this (according to Table 8 in HOLMES). In contrast, Sleuth [10] can detect suspicious behavior in specific steps through initial label propagation, but its label propagation mechanism may fail if the attacker uses legitimate tools (such as the command line) to operate, and cannot accurately identify malicious intent. Based on the context information of entities, these policies automatically determine whether an event is attack-related. The subject represents a process, the object represents different types of entities connected with the subject, and the direction indicates the information flow between subject and object. For example, the third policy represents process $\rightarrow$ file: if the process in the write event is considered attack-related, suspicious semantics will propagate from the process to the file. If another process reads the file containing suspicious semantics, the suspicious semantics will further propagate from the file to that process. Then, we use a heuristic search algorithm integrating equivalent semantic transitivity in Section IV-D.
TABLE III: Equivalent Semantic Transitivity Policies in the Context of Generalized Attack Pattern Identification and Matching.
@@
Event_Clone(p, p’)]: p’.semantics.add(”SSuspiciousLabel”)
@@
IV-D Threat Hunting and Incremental Aligning Module
IV-D1 Suspicious Semantic Tree Construction
@@
An event contains the interaction information between entities and can be transformed into an information flow, which can be further classified into data flows and control flows. Data flows indicate dependencies in data content, reflecting the data propagation path (e.g., a process reading a file), while control flows primarily refer to process creation relationships (e.g., a parent process creating a child process). In the threat hunting module, data flows and control flows will be jointly abstracted into a suspicious semantic tree. The process of generating a suspicious semantic tree is detailed as following three steps, as shown in algorithm 1:
@@
Step 1: Finding Candidate Nodes. To capture malicious behaviors in the provenance graph constructed from low-level system logs that match the patterns in the corresponding query graph, our system first searches for all nodes in the provenance graph with attributes identical to those of entity node in the query graph. These candidate nodes are collected into a list, referred to as the candidate set FC(i), which is associated with the query node (Line 7).
@@
Step 2: Confirming Attack Intent. As the query graph $G_{q}$ carries clear temporal features and causal relationships, we leverage these information to guide the attack detection reconstruction and reconstruction processes. This enables us to quickly determine the initial intrusion location, relevant entity nodes, and the sequence of attack events. Such a query graph can assist analysts in searching malicious behaviors effectively. The ExtractRelevantNodes function is subsequently invoked to identify the critical nodes within the query graph that are essential for comprehending the attack methodology (Line 8). This function operates by determining the next potential action nodes in $G_{p}$ based on the preceding step $F_{c}$ in $G_{q}$. The ReconstructAttackSequence algorithm reconstructs the attack sequence (Line 9), carefully aligning with the temporal and causal patterns within $G_{q}$. Unlike indiscriminate search methods, the approach is strategically guided by specific target nodes, precisely capturing the attacker’s intended progression. The function determines the matching order by meticulously tracing the sequence of attacks delineated in the graph. Actminer evaluates whether nodes introduce malicious semantics, which is identified by first fixing the target nodes and then analyzing behavior within $G_{q}$. This process involves tracing actions from the fixed nodes to determine patterns indicating malicious intent, forming the basis for further analysis.
@@
To optimize the traversal and analysis process, the algorithm stores context information for each node (Lines 11-15). This step avoids redundant traversals of the graph and enables the algorithm to order the edges based on timestamps, resulting in an ordered hunting sequence (Lines 16-19). This sequence is instrumental in guiding the subsequent steps of the algorithm. It will calculate the reciprocal of the length of the shortest path between nodes in Gq and Gp as the path score. For each node in Gq, the candidate node with the highest contribution value will be selected as the fixed node.
@@
Step 3: Creating Tree Nodes. The final step of the algorithm involves the creation of tree nodes (Lines 19-27). Given the disparity in size between the query graph and the provenance graph, the CreateTreeNode function (Line 31) is utilized to map each query graph to one or more subgraphs in the provenance graph that exhibit similar patterns. The function specifically creates a branch in the tree for the attack entry point. This mapping is facilitated by the hunting sequence obtained in the previous step, enabling efficient detection navigation against the large and complex dataset.
@@
After creating the tree nodes, when the attack progression exceeds the total attack sequence, the system raises an alert to notify security analysts of this anomalous situation.
@@
4:$G_{q}\leftarrow MergeSimilarEntity(G_{q})$
5:$L_{Se}\leftarrow GetQuerySequences(G_{q})$
6:$f\leftarrow GetSetQueryGraphFlow(G_{q})$
7:$G_{p}\leftarrow MergeSimilarEvent(G_{p})$
@@
9:$FC\leftarrow FindCandidateNodes(G_{q},G_{p})$
10:$relevant\_nodes\leftarrow ExtractRelevantNodes(G_{q})$
11:$attack\_sequence\leftarrow ReconstructAttackSequence(G_{q})$
12:$hunting\_sequence\leftarrow\emptyset$
@@
15:     $hunting\_sequence.add(context)$
@@
18:$visited\leftarrow\emptyset$
19:$fixed\_candidates\leftarrow\emptyset$
20:$seqNum\leftarrow 0$
@@
To illustrate the aforementioned methodology, we present an example for better comprehension. We aim to find suspicious subgraphs similar to the query graph in the provenance graph shown in Figure 4. As observed, the representation above illustrates a concrete instantiation of the graph, whereas the depiction below presents an abstraction of the model. Assuming we start from the process P1 within the red box as the starting point of the attack chain (corresponds to the above is $powershell_{1}.exe$), according to the query graph, the next step should be to find an executable file associated with P1, with the edge semantics being a write event. In the graph, only the Fe1 node (corresponds to the above is $update.ps1$) satisfies this condition, so we can generate a tree node that stores a variety of data, including the unique identifier of the query graph, event type, and relevant temporal information to assist subsequent hunting tasks. This step is largely analogous for both approaches, with only marginal differences in storage efficiency and temporal performance, the detail can see in Section V-D.
@@
IV-D2 Incremental Aligning
As time progresses, the log data generated by hosts will continue to increase. For traditional threat hunting systems, for any newly added logs after a period of time, they need to re-scan the entire dataset with a larger volume. To address the inefficiency of traditional threat hunting in analyzing incremental streaming data, we adopt an incremental graph computation method to hunt for attacks and update suspicious semantic trees. First, we search for new candidate nodes in the newly arrived provenance graph based on the attributes of nodes in the query graph. Next, we divide the impact of the new data on the suspicious subtrees into two parts: the new data affects the existing suspicious semantic subtrees, and the new data is unrelated to the existing suspicious subtrees. Furthermore, to manage memory overhead effectively, we implement a forgetting rate to reduce memory consumption. We transfer the nodes that remain un-updated for a period of 6 hours (can be adjusted according to different circumstances) into the database and create a corresponding index for them. The index includes the attributes of the node itself and its parent node, enabling rapid localization of relevant nodes in the event of a subsequent occurrence. This allows for the efficient retrieval and reinstatement of these nodes into memory.
For the former, through the affected tree nodes, we can obtain the current attack progress and their mapped nodes in the provenance graph and query graph. Then, through the sequence of suspicious candidate nodes, we can determine the next suspicious entity to hunt. Finally, we judge whether the suspicious states are met in the candidate node set of the suspicious entity, and if so, we construct a new tree node.
For the latter, we first determine whether the node has candidate nodes. If there are no candidate nodes, it means that this part of the data does not have an entry point for attacks, indicating that this part of the data is considered benign. However, if there are candidate nodes, we will start rebuilding the suspicious subtree from its candidate nodes.

### v2

@@ 5 System Design
5.1 System Overview
Data Preparation Module (§ 5.2). The attack query graphs are extracted from threat intelligence reports, and provenance graphs are constructed based on extensive underlying logs. Duplicate events and orphan nodes within the provenance graphs undergo filtering, which is a necessary and common practice in existing work inam2023sok .
Casual Relation and Semantic Processing Module (§ 5.3). When a new attack query graph or provenance graph is generated, Actminer will first categorize the entities into four classes. The provenance graph will then deliver to the next module while the attack query graph still needs to be processed. Next, Actminer merges analogous actions in the attack query graph. Finally, Actminer employs EST, which traces potentially overlooked attack chains by tracking malicious semantics, to identify suspicious actions in the next module.
Threat Hunting and Incremental Aligning Module (§ 5.4). Actminer will hunt attack-related scenario by chaining suspicious semantic nodes and generating suspicious semantic tree. As time progresses, batch data is continuously inputted into the Actminer, persistently updating our suspicious semantic trees and unveiling more latent malicious attacks. To control the memory consumption, Actminer will store the unupdated tree branch to the database unless a certain behavior is observed related to this branch.
The basic architecture of Actminer is shown in Figure 2, which can be divided into three modules: (I) the Data Preparation Module, (II) Casual Relation and Semantic Processing Module, and (III) Threat Hunting and Incremental Aligning Module. It is important to note that Actminer continuously runs the above three modules as time progresses. Details of system design for each module are provided in Section 5.2, Section 5.3, and Section 5.4, respectively.
@@ 5.1 System Overview
5.2 Data Preparation Module
@@
5.2.1 Provenance Preparation
Provenance graphs are composed of log data collected from various platforms by data collectors. In this work, we employ open-source tools, such as eAuditd sekar2023eaudit , Kellect chen2022kellect , and Event Tracing for Windows etw , to gather relevant system logs from Linux and Windows environments. Actminer transforms each event into a directed, time-stamped edge, in which the source node represents the object being acted upon. For any event $et\in E$, Actminer represents it as a quintuple $\langle{UID_{s},UID_{o},OP,Ti_{i}}\rangle$. $UID_{s}$ and $UID_{o}$ are unique identifiers for the subject and object of $et$, respectively. $OP$ denotes the type of $et$, and $Ti_{i}$ denotes the time when $et$ occurred.
Directly processing such massive raw log data is extremely challenging. To address this, we perform pruning operations on the low-level log data. Specifically, redundant events without context are removed zhu2021general ; zhu2023aptshield . This means that if the subject UID ($UID_{s}$), object UID ($UID_{o}$), and operation ($OP$) are identical and the timestamps ($T_{i}$) are consecutive, the most recent timestamp will be preserved. Furthermore, our methodology includes the removal of isolated nodes within the provenance graph. The isolated nodes refer to entities that lack any incoming or outgoing edges. For example, we find that the data contains many of these nodes, which manifest without any parent or child nodes. Moreover, none of the events contain a subject UID ($UID_{s}$) or object UID ($UID_{o}$) matching the UID of the node, lack contextual information, and fail to provide meaningful insights. Therefore, they must be removed without compromising the integrity of the graph representation.
Concurrently, during the process of constructing provenance graphs, the node types are categorized based on the entity type contained within the logs. For instance, $F_{a}$ represents files involving user and system-sensitive information, such as the boot.ini file on Windows and /etc/passwd file on Linux. The specific categorization is summarized in Table 1.
As presented in Table 1, we are inspired by Aptshield zhu2023aptshield and Conan xiong2020conan . We then refined the classification heuristics and optimized the classification method we obtained from POIROT. Furthermore, we extensively gather CTI reports from various online sources and network channels, identifying specific file paths that exhibit heightened susceptibility to attacks. Consequently, we adapt the importance degree of these paths based on their frequency of occurrence to finally obtain ten distinct labels.
Table 1: A categorization of distinct entities and their corresponding label assignments.
@@ 5.2.1 Provenance Preparation
By analyzing CDM18 and CDM19, which refer to data definitions for DARPA’s E3 and E4 programs, respectively, we adopt events based on a few general fields in the CDM (i.e., events of read, write, fork, clone, create, execute, load, and inject), which were most commonly used in previous studies, such as Holme milajerdi2019holmes , Sleuth hossain2017sleuth , and MORSE hossain2020combating . The details are presented in Table 2.
Table 2: Classification of different event types.
@@
5.2.2 Query Graph Preparation.
@@
CTI reports describe attacks that have already occurred. We collect the latest threat intelligence from websites such as Microsoft and Symantec. Leveraging open-source tools such as Extractor satvat2021extractor , we extract attack query graphs ($G_{q}$) from CTI reports. In our extraction process, we found that the graphs extracted by the Extractor exhibit issues, such as incompleteness and fragmentation. CRUcialG cheng2024crucialg was then developed specifically to address and solve this issue. The graphs extracted by the Extractor are input into CRUcialG, and the graphs can represent the attack process and semantics by combining automation with expert experience.
We construct query graphs by analyzing CTI attack reports and abstracting attack patterns into graph representations. CTI reports are compiled by security analysts to document IOCs and malicious activities. Figure 3 illustrates an example query graph constructed from an attack report in the DARPA TC3 dataset darpa-tc . Processes, files, and network flows are represented by rectangles, ovals, and diamonds, respectively. The attack initiates by compromising a Firefox browser extension, subsequently downloading a malicious executable file (“ztmp”) to the target host and establish initial access. The attack then executes this malicious file, creating a corresponding malicious process (“ztmp”) that connects to a command-and-control (C&C) server via the external IP address 162.66.239.75. The malicious process conducts internal reconnaissance by scanning for open ports using “portscan” before deleting the malicious file to eliminate attack artifacts. Figure 4 shows the query graph after processing.
Figure 3: The type and details of query graph
Figure 4: The type and details of the processed query graph
5.3 Casual Relation and Semantic Processing Module
@@
In this section, we sequentially describe the process of Module II in Figure 2, i.e., merging analogous actions and EST. Briefly, analogous actions in attack query graphs are merged while employing an EST strategy. This process enhances attack hunting capabilities while simultaneously preparing the data for input to Module III.
5.3.1 Merging Analogous Actions
@@
When Actminer adopts the attack query graphs to perform the threat hunting tasks, it encounters the issue of rigid matching, causing the difficulty to achieve alignment between attack query graphs and provenance graphs. Finally, it would result in high FN and FP rate in Module III.
@@
Additionally, the directly extracted $G_{q}$ is fixed, limiting its ability to generalize and effectively defend against variants of known attacks, as it is specifically tailored to documented attack patterns. A key insight of Actminer, merging analogous actions, is developed to address this challenge. This method integrates nodes with similar operations. For instance, instead of representing each individual file with a distinct node, the generalized $G_{q}$ may group files based on their types (e.g., $F_{a},F_{b},F_{c}$ in Table 1) in the attack chain. Simultaneously, we consider the temporal relationships of attack events. Processes can be merged based on their functional attributes or the operations they perform, rather than being strictly tied to specific executable paths or process names. In essence, this method transforms the information flow from the source entity to the same target entity while preserving the semantic meaning of the source entity, so that equivalent events can be removed as redundant information, thereby achieving efficient compression of the data.
5.3.2 Equivalent Semantic Transfer
@@
During the attack process, malicious processes controlled by the attacker interact with other entities xiong2020conan , causing malicious behaviors and effects to proliferate across the system through the intricate web of entity interactions and information propagation pathways and expanding the attacker’s control scope.
@@
Based on above finding, we construct the equivalent semantic transitivity strategy to address the FN issue. The key insight of this strategy is that the semantics of malicious behavior propagate with taking actions. Figure 5.(a) depicts a short attack path described in the query graph, where a process controlled by the attacker tampers with sensitive files. Figure 5.(b) illustrates one of the attacker’s specific implementation approaches. First, a process P1 with suspicious semantics injects malicious code to create process P2, and then process P2 writes a suspicious file Fe. Next, another malicious process P3 accesses the tainted file. Subsequently, process P3 with suspicious semantics creates process P4. Finally, process P4 with suspicious semantics performs a write operation on sensitive files. Through this series of operations, the ultimate behavior of a suspicious process tampering with sensitive files Fa is achieved. Although the types of controlled entities change during the attack process, the ultimate goal remains unchanged. The operations in (b) are essentially the same as those in (a) and align with the attacker’s intent, making the two paths semantically equivalent.
@@
To comprehensively capture attack intents and mitigate attack evasion while addressing semantic gaps between $G_{q}$ and $G_{p}$, we integrated log data from the DARPA project, leveraged attack stage theories proposed in previous works such as Conan xiong2020conan and Aptshield zhu2023aptshield , and incorporated relevant descriptions from the CDM document DARPA3-CDM to extract six equivalent semantic transitivity policies, as presented in Table 3. Therefore, Actminer can track and locate suspicious behavior on the host in real time by analyzing the data flow transmission, while systems like HOLMES milajerdi2019holmes cannot do this (according to Table 8 in HOLMES). In contrast, Sleuth hossain2017sleuth employs initial label propagation to detect suspicious behavior in specific steps. For regular processes, a malicious command line affects only the data trustworthiness tag (data t-tag), leaving the code trustworthiness tag (code t-tag) as “benign.” As a result, detection policies focused on the code t-tag do not trigger alerts based solely on the command line itself. Based on the context information of entities, these policies automatically determine whether an event is attack related or not. The subject represents a process, the object represents different types of entities connected with the subject, and the direction indicates the information flow between the subject and the object. For example, the third policy represents process $\rightarrow$ file: if the process in the write event is considered as attack related, suspicious semantics will propagate from the process to the file. If another process reads the file containing suspicious semantics, the suspicious semantics will further propagate from the file to that process. We then use a heuristic search algorithm integrating equivalent semantic transitivity in Section 5.4.
Table 3: Equivalent Semantic Transitivity Policies in the Context of Generalized Attack Pattern Identification and Matching.
@@
Event_Clone(p, p’)]: p’.semantics.add("SSuspiciousLabel")
@@
5.4 Threat Hunting and Incremental Aligning Module
5.4.1 Suspicious Semantic Tree Construction
@@
An event contains the interaction information between entities and can be transformed into an information flow, which can be further classified into data flows and control flows. Data flows indicate dependencies in data content, reflecting the data propagation path (e.g., a process reading a file), while control flows primarily refer to process creation relationships (e.g., a parent process creating a child process). In the threat hunting module, data flows and control flows will be jointly abstracted into a suspicious semantic tree. The following shows the process of generating a suspicious semantic tree, as presented in algorithm 1:
@@
Step 1: Finding Candidate Nodes. To capture malicious behaviors in the provenance graph constructed from low-level system logs that match the patterns in the corresponding query graph, our system first searches for all nodes in the provenance graph with attributes identical to those of the entity node in the query graph. These candidate nodes are collected into a list, referred to as the candidate set FC(i), which is associated with the query node (line 7).
@@
Step 2: Confirming Attack Intent. As the query graph $G_{q}$ carries clear temporal features and causal relationships, we leverage these information to guide the attack detection reconstruction and reconstruction processes. This enables us to quickly determine the initial intrusion location, relevant entity nodes, and sequence of attack events. Such a query graph can assist analysts in searching malicious behaviors effectively. The ExtractRelevantNodes function is subsequently invoked to identify the critical nodes within the query graph that are essential for comprehending the attack methodology (line 8). This function operates by determining the next potential action nodes in $G_{p}$ based on the preceding step $F_{c}$ in $G_{q}$. The ReconstructAttackSequence algorithm reconstructs the attack sequence (line 9), carefully aligning with the temporal and causal patterns within $G_{q}$. Unlike indiscriminate search methods, the approach is strategically guided by specific target nodes, accurately capturing the attacker’s intended progression. The function determines the matching order by meticulously tracing the sequence of attacks delineated in the graph. Actminer evaluates whether nodes introduce malicious semantics, which is identified by first fixing the target nodes and then analyzing behavior within $G_{q}$. This process involves tracing actions from the fixed nodes to determine patterns indicating malicious intent, forming the basis for further analysis.
@@
To optimize the traversal and analysis process, the algorithm stores context information for each node (lines 11–15). This step avoids redundant traversals of the graph and enables the algorithm to order the edges based on timestamps, resulting in an ordered hunting sequence (lines 16–19). This sequence is instrumental in guiding the subsequent steps of the algorithm. It will calculate the reciprocal of the length of the shortest path between nodes in Gq and Gp as the path score. For each node in Gq, the candidate node with the highest contribution value will be selected as the fixed node.
@@
Step 3: Creating Tree Nodes. The final step involves the creation of tree nodes (lines 19–27). Given the disparity in size between the query graph and the provenance graph, the CreateTreeNode function (line 31) is used to map each query graph to one or more subgraphs in the provenance graph that exhibit similar patterns. The function specifically creates a branch in the tree for the attack entry point. This mapping is facilitated by the hunting sequence obtained in the previous step, enabling efficient detection navigation against the large and complex dataset.
@@
After the tree nodes are created, when the attack progression exceeds the total attack sequence, the system raises an alert to notify security analysts of this anomalous situation.
@@
[…]
@@
[…]
@@
15:  hunting_sequence.add(context)hunting\_sequence.add(context)
@@
17:hunting_sequence.sort(key=lambdax:x.timestamp)hunting\_sequence.sort(key=lambdax:x.timestamp)
@@
[…]
@@
We present an example for better comprehension to illustrate the aforementioned methodology. We aim to find suspicious subgraphs similar to the query graph in the provenance graph shown in Figure 6. The representation above illustrates a concrete instantiation of the graph, whereas the depiction below presents an abstraction of the model. Assuming we start from process P1 within the red box as the starting point of the attack chain (corresponding to the above is $powershell_{1}.exe$), according to the query graph, the next step should be to find an executable file associated with P1, with the edge semantics being a write event. In the graph, only the Fe1 node (corresponding to the above is $update.ps1$) satisfies this condition; therefore, we can generate a tree node that stores a variety of data, including the unique identifier of the query graph, event type, and relevant temporal information to assist subsequent hunting tasks. This step is largely analogous for both approaches, with only marginal differences in storage efficiency and temporal performance. The detail is provided in Section 6.4.
@@
5.4.2 Incremental Aligning
As time progresses, the log data generated by hosts will continue to increase. For any newly added logs after a period of time, traditional threat hunting systems need to rescan the entire dataset with a larger volume. To address the inefficiency of traditional threat hunting systems in analyzing incremental streaming data, we adopt an incremental graph computation method to hunt for attacks and update suspicious semantic trees. First, we search for new candidate nodes in the newly arrived provenance graph based on the node attributes in the query graph. Next, we divide the impact of new data on the suspicious subtrees into two parts, i.e., new data that affects the existing suspicious semantic subtrees and new data that is unrelated to the existing suspicious subtrees. Furthermore, to manage memory overhead effectively, we implement a forgetting rate to reduce memory consumption. We transfer the nodes that remain unupdated for a period of 6 h (this duration can be adjusted according to different circumstances) into the database and create a corresponding index for them. The index includes the attributes of the node itself and its parent node, enabling rapid localization of relevant nodes in the event of a subsequent occurrence. This allows for the efficient retrieval and reinstatement of these nodes into memory.
For the former, through the affected tree nodes, we can obtain the current attack progress and their mapped nodes in the provenance graph and query graph. Through the sequence of suspicious candidate nodes, we can then determine the next suspicious entity to hunt. Finally, we judge whether the suspicious states are met in the candidate node set of the suspicious entity. If this is achieved, we then construct a new tree node.
For the latter, we first determine whether the node has candidate nodes. The absence of candidate nodes means that this part of the data does not have an entry point for attacks, indicating that this part of the data is considered benign. However, if there are candidate nodes, we will start rebuilding the suspicious subtree from its candidate nodes.